# The causal structure of galactic astrophysics

### v2

@@ 1 Introduction
Understanding the physical processes that shape galaxies is a central goal of astrophysics. Empirical progress has traditionally relied on identifying correlations between observed properties, which can then be interpreted in light of theoretical models for galaxy formation and used to constrain them. The advent of large surveys and powerful machine learning techniques has greatly expanded our ability to find such statistical associations, uncovering intricate patterns across high-dimensional parameter spaces. However, correlation alone cannot determine causal influences among variables: which properties are actually responsible for determining others, in what direction this influence goes, and whether there exist confounding variables that are not included in the dataset but influence those that are. Achieving this requires causal discovery, a methodology widely applied in fields such as genomics, epidemiology and economics, but that has had extremely limited exposure in the physical sciences.
@@
A possibly complete list of previous applications of causal discovery to astrophysics follows. Mucesh et al. (2024) estimate a causal model of galaxy formation from semi-analytic models and hydrodynamical simulations, and compare it to non-causal models. Pasquato et al. (2023) apply the Peter–Clark and Fast Causal Inference algorithms to 83 galaxies in an attempt to constrain evolution mechanisms for their central supermassive black holes. Jin et al. (2024); Jin et al. (2025b) addresses the same question with 101 galaxies, using a Bayesian method for estimating the probabilities of all possible causal structures. Davis et al. (2025) use Fast Causal Inference to help constrain the formation histories of trans-Neptunian objects. Finally, Jin et al. (2025a) apply a linear causal discovery model to 100 simulated galaxies to constrain chemodynamical pathways relevant for the Milky Way. These studies involve samples too small to be representative of the underlying populations, and have been restricted to very specific variable sets and scientific questions.
After describing causal discovery and identifying a method suitable for astrophysical problems involving hundreds of thousands of objects, we illustrate the technique with galaxy data. Specifically we take seven variables describing the first-order properties (brightness, mass, size, morphology, star formation rate) of $\sim 5\times 10^{5}$ low-redshift galaxies from the Nasa Sloan Atlas (NSA). We assess reliability of the causal discovery and calibrate hyperparameters of our algorithm on mock data, then infer the causal links present in the NSA including the presence of confounding variables not in the data subset. We show explicitly how this can be used to refine our understanding of galaxy evolution and test proposed physical mechanisms, which crucially (as is typical in physics) come with a direction of operation in addition to simply inducing correlation among galaxy properties.
@@ 2 Observational data
  • elpetro_b300: current star formation rate (SFR) divided by the average over the past 300 Myr.
@@
These are designed to capture the most important information about galaxy structure, including mass, luminosity, size, structure and redshift. Note that they are by no means unique: different choices would reflect different user priorities and beliefs about the most causally relevant variables. The quantities designated ‘elpetro’ derive from elliptical Petrosian flux fits, which are deemed the most reliable in the catalogue. elpetro_mass and elpetro_b300 have been $K$-corrected to rest-frame magnitudes using the kcorrect code (Blanton & Roweis, 2007). Absolute magnitudes are given with $H_{0}=100h$ km/s/Mpc so should be read as $M-5\log{h}$. All logarithms are base 10.
To clean the catalogue we require zdist <0.15<0.15, elpetro_absmag <10<-10, elpetro_b300 $10^{-8}$, elpetro_b300 <10<10, elpetro_mass $10^{6}$, elpetro_mass $10^{12}$, sersic_n <6<6, elpetro_ba >0>0, elpetro_ba <1<1, elpetro_th50_r >0>0 and elpetro_th50_r <25<25. These cuts remove anomalous objects whose properties are likely to have been inaccurately determined. This leaves 587,338 out of an original 641,409 galaxies. A corner plot of the final dataset is shown in Fig. 1.
@@ 3.1 Causal Discovery
Most statistical analyses in astrophysics (whether or not through the lens of machine learning) are designed to measure correlations: how strongly two quantities co-vary and the properties of their relationship. Correlation, however, is agnostic about direction and mechanism, which are the predictions of galaxy formation theories and hence the most useful features for testing them. Projecting these predictions onto the space of correlations loses information and hence constraining power. Causal discovery methods seek to retain this information, going beyond correlation by inferring the causal structure that generates the observed data. This provides added value to the results that complements or is overlaid upon the traditional results of astrostatistical methodology. It is this information that connects directly to theories and models of galaxy formation and evolution. For thorough reviews of causal discovery see Spirtes et al. (1993); Mooij et al. (2014); Eberhardt (2017); Glymour et al. (2019).
@@
Causal discovery utilises conditional correlation strengths to uncover the directions of influence among variables. To see how this works, suppose we measure three galaxy properties: stellar mass ($M$), gas mass ($G$), and star-formation rate ($S$). We find all three are correlated: larger values of one are associated with larger values of any other. From this alone, one could imagine a mass-driven scenario (more massive galaxies accrete more gas, $M\rightarrow G$, which in turn fuels star formation, $G\rightarrow S$), a feedback-driven scenario (high $S$ regulates gas supply, $S\rightarrow G$, while simultaneously building stellar mass, $S\rightarrow M$) or a common-cause scenario (the environment controls both mass growth and gas supply, indirectly correlating all three). However, if $M$ and $S$ remain correlated even after conditioning on $G$, there must exist a direct causal link between them ($M\rightarrow S$), while if conditioning on $G$ removes the MSM-S correlation we must instead have $M\rightarrow G\rightarrow S$. Thus the existence or absence of the conditional correlation breaks the degeneracy between the physical scenarios.
@@
By assessing all such conditional correlations (including a multi-dimensional conditioning set) one can determine the Markov equivalence class to which the data belongs. Each class comprises causal structures sharing the same statistical dependencies which therefore cannot be distinguished without experimental intervention, impossible in astrophysics. (Such statistical independencies can also be thought of as implying factorisability of the joint (probability) distribution describing the variables: e.g. $A\rightarrow B\rightarrow C$ implies $P(A,B,C)=P(A)P(B|A)P(C|B)$.) This leads to the classic constraint-based causal discovery method, the Peter–Clark algorithm, which eliminates direct (i.e. causal) correlations with conditional independence tests then applies orientation rules to fill in directions where possible.  [^2] As an example of such a rule, suppose that one has identified the direct links ABA-B, ACA-C, BDB-D, CDC-D. If $B$ and $C$ are disconnected when conditioning on $A$ but not when conditioning on $D$, it must be that neither $B$ nor $C$ are caused by $D$. Hence the BDB-D and CDC-D links must be $B\rightarrow D$ and $C\rightarrow D$. There are two more rules involving collisions that can be used to determine directions. Alternative score-based methods such Greedy Equivalent Search assign likelihoods to candidate solutions based on the correlation strengths and search for the highest-scoring structures, while functional causal models such as additive noise models instead exploit asymmetries in functional relationships to determine causal direction.
Causal structure is visualised in a causal graph—a directed network of relationships among the variables. The true, generating causal structure is described by a directed acyclic graph (DAG), in which all edges (i.e. causal links between variables, represented by lines) are directed. The corresponding Markov equivalence class is shown by a completed partially directed acyclic graph (CPDAG), which contains edges for which directionality cannot be established based solely on the conditional independencies. A partial ancestral graph (PAG) additionally models the potential effect of latent confounding variables, overcoming the assumption that all causes are contained within the dataset. These can create dependencies between observables that cannot be resolved by any orientation of arrows among the observed variables alone: every candidate orientation produces an inconsistency with other independencies. This replaces the DAG with a maximal ancestral graph (MAG). A PAG then represents the Markov equivalence class of the true MAG, indicating the possible influence of hidden causes with a circle endpoint. This is the goal of causal discovery.
@@
Several assumptions are required for causal discovery to be possible. The most common are the Markov condition (separated variables in the causal graph are statistically independent), faithfulness (no accidental statistical independences) and acyclicity (nothing can be indirectly its own cause). In addition, the conditional correlations must be adequately captured by the statistical test applied (which come with their own assumptions) and the threshold $p$-value chosen to identify insignificant correlations. Some methods further assume causal sufficiency, namely that all causally relevant variables are included in the dataset. Since causal discovery cannot guarantee the “true” causal graph, but rather identifies the set of structures compatible with the observed data and assumptions, it is best used to weed out statistically implausible causal relationships and hence generate testable causal hypotheses to be compared with theoretical expectations or followed up with targeted observations.
@@ 3.2 The FCIT algorithm applied to galaxies
For application to the NSA, we have the requirements that a method is 1) accurate in the presence of confounding latent variables, since it is highly unlikely that all relevant information is contained in the dataset, 2) able to accommodate highly non-linear functional relationships (see Fig. 1), and 3) efficient enough to analyse $N\approx\mathcal{O}(10^{5})$ objects in reasonable time. This is impossible for traditional algorithms, which scale as $N^{3}$ or worse when accounting for confounders and nonlinearity and become prohibitively memory intensive with $\gtrsim 500$ samples (Strobl et al., 2017; Raghu et al., 2018).
To achieve this we adopt the newly-developed method Fast Causal Inference with Targeted Testing (FCIT; Ramsey et al. 2025a) as implemented in py-tetrad (Ramsey & Andrews, 2023). [^3] https://www.cmu.edu/dietrich/philosophy/tetrad/use-tetrad/tetrad-python.html, [^4] https://github.com/cmu-phil/py-tetrad This is a hybrid constraint-and-score-based algorithm which starts with a score-based estimate of the causal graph, possibly with additional adjacencies that the score-based method cannot remove directly, and then applies a recursive procedure to identify minimal sets of conditional independence tests for removing these extra causal links between variables that should be non-adjacent in the correct PAG. This reduces the number of tests required and the statistical noise they introduce. It also incorporates discriminating path checks during edge removal, ensuring that edges are properly oriented before deciding on conditional independence. Additional bookkeeping in the FCIT procedure guarantees that interim and final graphs are guaranteed to be well-formed PAGs when applied to data samples, a departure from existing latent variable procedures that yield PAGs theoretically. The resulting graphs are edge-minimal, correctly oriented and exhibit high accuracy on causal discovery benchmarks, with an unprecedented runtime of only $\sim$1 minute for $5\times 10^{5}$ datapoints and 7 features.
For independence test we use use_basis_function_lrt, which expands variables into a nonlinear basis set and performs a likelihood ratio test between different conditional independence models. For scoring we adopt use_basis_function_bic, which uses the same nonlinear expansion but employs a Bayesian information criterion (BIC) score with a custom penalty weight:
@@
where $\mathcal{L}$ is the likelihood of a target variable according to a hypothetical conditional correlation structure, $N$ is the sample size used for the local regression, and $k$ is the number of free regression coefficients in the basis‑function expansion for the child given its current parent set (Ramsey et al., 2025b). This makes FCIT a hybrid algorithm: the constraint phase uses the basis‑function likelihood ratio test to prune edges, while the score phase (and the refinement and orientation phases) uses the basis‑function BIC. For the $p$-value threshold we use 0.01 in all cases, which we find not to affect the results appreciably.
Two main hyperparameters affect performance in a dataset-specific way: penalty_discount and truncation_limit. The former controls how strongly graph complexity is penalised (Eq. 1). A higher value favours simpler graphs by removing more noise-dominated edges, at the expense of the quality of the conditional fit as described by $\mathcal{L}$. The latter controls the complexity of the local regression model through the number of polynomial basis terms included. Larger values allow more expressive models at the cost of runtime and enhanced BIC complexity penalty.

### v3

@@ 1 Introduction
Understanding the physical processes that shape galaxies is a central goal of astrophysics. Empirical progress has traditionally relied on identifying correlations between observed properties, which can then be interpreted in light of theoretical models for galaxy formation and used to constrain them. The advent of large surveys and powerful machine learning techniques has greatly expanded our ability to find such statistical associations, uncovering intricate patterns across high-dimensional parameter spaces. However, correlation alone cannot determine causal influences among variables: which properties are actually responsible for determining others, in what direction this influence goes, and whether there exist confounding variables that are not included in the dataset but influence those that are (Pearl, 2009). Achieving this requires causal discovery, a methodology widely applied in fields such as genomics, epidemiology and economics, but that has had extremely limited exposure in the physical sciences (Mooij et al., 2016; Glymour et al., 2019a).
@@
A possibly complete list of previous applications of causal discovery to astrophysics follows. Pasquato et al. (2023) apply the Peter–Clark and Fast Causal Inference algorithms to 83 galaxies in an attempt to constrain evolution mechanisms for their central supermassive black holes. Jin et al. (2024, 2025b) addresses the same question with 101 galaxies, using a Bayesian method for estimating the probabilities of all possible causal structures. Mucesh et al. (2024) estimate a causal model of galaxy formation from semi-analytic models and hydrodynamical simulations, and compare it to non-causal models. (They focus mainly however on causal inference—the estimation of causal effects given a causal model—rather than causal discovery, the learning of the causal model in the first place.) Davis et al. (2025) use Fast Causal Inference to help constrain the formation histories of trans-Neptunian objects. Zhang et al. (2025) apply a Fast Causal Inference-like algorithm to learn the interrelations between photometric properties of $\sim$200 galaxy clusters. Finally, Jin et al. (2025a) apply a linear causal discovery model with small conditioning set size and the assumption of Gaussian residuals to $\sim 150000$ stars in simulated galaxies to constrain chemodynamical pathways relevant for the Milky Way. These studies involve samples either too small to be representative of the underlying populations, or, in the latter case, making stringent assumptions about the correlation structure that may not hold in real astrophysical data.
After describing causal discovery and identifying a method suitable for astrophysical problems involving hundreds of thousands of objects, we illustrate the technique with galaxy data. Specifically we take seven variables describing the first-order properties (brightness, mass, size, morphology, star formation rate) of $\sim 4.5\times 10^{5}$ low-redshift galaxies from the Nasa Sloan Atlas (NSA). We assess reliability of the causal discovery and calibrate hyperparameters of our algorithm on mock data, then infer the causal links present in the NSA including the presence of confounding variables not in the data subset. We show explicitly how this can be used to refine our understanding of galaxy evolution and test proposed physical mechanisms, which crucially (as is typical in physics) come with a direction of operation in addition to simply inducing correlation among galaxy properties.
@@ 2 Observational data
  • elpetro_b300: current star formation rate (SFR) divided by the average over the past 300 Myr. This indicates how recently a galaxy formed most of its stars since 300 Myr ago.
@@
These are designed to capture the most important information about galaxy structure, including mass, luminosity, size, structure and redshift. Note that they are by no means unique: different choices would reflect different user priorities and beliefs about the most causally relevant variables. The quantities designated ‘elpetro’ derive from elliptical Petrosian flux fits, which are deemed the most reliable in the catalogue. elpetro_mass and elpetro_b300 have been $K$-corrected to rest-frame magnitudes using the kcorrect code (Blanton and Roweis, 2007). Absolute magnitudes are given with $H_{0}=100h$ km/s/Mpc so should be read as $M-5\log{h}$. All logarithms are base 10.
To clean the catalogue we require zdist <0.15<0.15, elpetro_absmag <10<-10, elpetro_b300 $10^{-8}$, elpetro_b300 <10<10, elpetro_mass $10^{6}$, elpetro_mass $10^{12}$, sersic_n <5.8<5.8, elpetro_ba >0>0, elpetro_ba <1<1, elpetro_th50_r >0>0 and elpetro_th50_r <25<25. These cuts remove anomalous objects whose properties are likely to have been inaccurately determined. This leaves 445,763 out of an original 641,409 galaxies. A corner plot of the final dataset is shown in Fig. 1.
@@ 3.1 Causal Discovery
Most statistical analyses in astrophysics (whether or not through the lens of machine learning) are designed to measure correlations: how strongly two quantities co-vary and the properties of their relationship. Correlation, however, is agnostic about direction and mechanism, which are the predictions of galaxy formation theories and hence the most useful features for testing them. Projecting these predictions onto the space of correlations loses information and hence constraining power. Causal discovery methods seek to retain this information, going beyond correlation by inferring the causal structure that generates the observed data. This provides added value to the results that complements or is overlaid upon the traditional results of astrostatistical methodology. It is this information that connects directly to theories and models of galaxy formation and evolution. For thorough reviews of causal discovery see Spirtes et al. (1993); Mooij et al. (2016); Eberhardt (2017); Glymour et al. (2019b).
@@
Causal discovery utilises conditional correlation strengths to uncover the directions of influence among variables. To see how this works, suppose we measure three galaxy properties: total dynamical mass enclosed within the extent of the galaxy ($M$, mainly driven by dark matter), gas mass ($G$), and star-formation rate ($S$). We find all three are correlated: larger values of one are associated with larger values of any other. From this alone, one could imagine a mass-driven scenario (more massive galaxies accrete more gas, $M\rightarrow G$, which in turn fuels star formation, $G\rightarrow S$), a feedback-driven scenario (high $S$ regulates gas supply, $S\rightarrow G$, while simultaneously drawing dark matter in through adiabatic contraction and hence increasing $M$, $S\rightarrow M$Blumenthal et al. 1986) or a common-cause scenario (the environment controls both mass growth and gas supply, indirectly correlating all three). However, if $M$ and $S$ are decorrelated by conditioning on $G$ then $G$ must be responsible for linking them, which in the context of the above trichotomy implies the mass-driven scenario. Conversely, if $M$ and $S$ remain correlated even after conditioning on $G$, there must either exist a direct causal link between them (the feedback-driven scenario), or a latent variable links them irrespective of $G$ (the common-cause scenario). This is illustrated in Fig. 2. Thus the existence or absence of the conditional correlation breaks the degeneracy between the physical scenarios.
Figure 2: A visual representation of the example of Sec. 3.1 illustrating how causal discovery can distinguish between physical mechanisms beyond mere correlations. The axes of the scatter plots show the residuals of $S$ and $M$ upon subtraction of their expectation given their $G$ values and the average GSG-S and GMG-M relations. The upper causal graph has a conditional independence between $S$ and $M$ at fixed $G$, while the lower two graphs do not—the presence or absence of this conditional independence therefore narrows down the possible causal structure of the system.
@@
By assessing all such conditional correlations (including a multi-dimensional conditioning set) one can determine the Markov equivalence class to which the data belongs (Verma and Pearl, 1990). Each class comprises causal structures sharing the same statistical dependencies which therefore cannot be distinguished without experimental intervention, impossible in astrophysics. (Such statistical independencies can also be thought of as implying factorisability of the joint (probability) distribution describing the variables: e.g. $A\rightarrow B\rightarrow C$ implies $P(A,B,C)=P(A)P(B|A)P(C|B)$.) This leads to the classic constraint-based causal discovery method, the Peter–Clark algorithm (Spirtes and Glymour, 1991), which eliminates direct (i.e. causal) correlations with conditional independence tests then applies orientation rules to fill in directions where possible. [^2] As an example of such a rule, suppose that one has identified the direct links ABA-B, ACA-C, BDB-D, CDC-D. If $B$ and $C$ are disconnected when conditioning on $A$ but not when conditioning on $D$, it must be that neither $B$ nor $C$ are caused by $D$. Hence the BDB-D and CDC-D links must be $B\rightarrow D$ and $C\rightarrow D$. There are two more rules involving collisions that can be used to determine directions. Alternative score-based methods such Greedy Equivalence Search (Chickering, 2003) assign likelihoods to candidate solutions based on the correlation strengths and search for the highest-scoring structures, while functional causal models such as additive noise models instead exploit asymmetries in functional relationships to determine causal direction.
Causal structure is visualised using a causal graph, a graphical representation in which nodes correspond to observed variables and edges encode direct causal relationships. The true data-generating process is described by a directed acyclic graph (DAG), in which all edges are directed (represented by an arrow) and causal cycles are forbidden (Pearl, 2009). A DAG encodes both causal relations and a set of conditional independence constraints via the Markov property: two variables are independent given a conditioning set if all paths connecting them in the causal graph are blocked by the conditioned variables, where causal colliders along the path (two variables mutually causing a third) do not have conditioned descendants. This implies that there is no link between them through variables outside the conditioning set, an important constraint on their causal relation (Lauritzen, 1996). However, when only observational data are available the full DAG is typically not identifiable and only the Markov equivalence class can be determined. This is represented by a completed partially directed acyclic graph (CPDAG), in which some edges are oriented while others remain undirected, indicating causal directions that cannot be uniquely determined from conditional independencies alone (Andersson et al., 2000; Chickering, 2003).
The CPDAG representation relies on the assumption of causal sufficiency, namely that all common causes of the observed variables are themselves observed. When this assumption is violated, latent (i.e. unobserved) confounding variables may induce statistical dependencies that cannot be explained by any DAG over the observed variables alone (Spirtes et al., 1993). In such cases, every possible orientation of edges among observables leads to inconsistencies with the observed conditional independence structure. To accommodate this, the DAG over all variables (observed and latent) can be projected onto the observed variables, yielding a maximal ancestral graph (MAG; Richardson and Spirtes 2000). A MAG allows for both directed edges as in a DAG or CPDAG, representing direct causal influence, and bidirected edges ($\leftrightarrow$) , representing the presence of an unobserved common cause.
As with DAGs, MAGs are generally not uniquely identifiable from observational data because the conditional independence structure does not pick out a specific graph. The equivalence-class representation in this setting is the partial ancestral graph (PAG), which summarises all MAGs consistent with the observed conditional independences (Zhang, 2008). In a PAG, edge endpoints may be marked with arrowheads, tails, or circles. An arrowhead ($\rightarrow$) indicates that the endpoint cannot be a cause of the adjacent variable in any compatible MAG, while a tail (-) indicates that it must be an cause. A circle endpoint ($\circ$) denotes ambiguity, meaning the corresponding ancestral relationship cannot be resolved and may be impacted by latent confounders. For example, an edge $X\circ-\circ Y$) implies consistency with a direct causal effect between $X$ and $Y$ in either direction, a bidirected edge due to an unobserved confounder, or a combination thereof. PAGs therefore encode both causal direction uncertainty and uncertainty arising from hidden variables. When analysing data using conditional independence information, recovering a PAG is the aim of causal discovery algorithms that do not assume causal sufficiency. This is typically crucial in real-world applications where one cannot be sure that all causally relevant variables are included in one’s dataset.
@@
Several assumptions are required for causal discovery to be possible. The most common are the Markov condition (separated variables in the causal graph are statistically independent), faithfulness (no accidental statistical independences) and acyclicity (nothing can be indirectly caused by itself). In addition, the conditional correlations must be adequately captured by the statistical test applied (which come with their own assumptions) and the threshold $p$-value chosen to identify insignificant correlations. Some methods further assume causal sufficiency, namely that all causally relevant variables are included in the dataset. Since causal discovery cannot guarantee the “true” causal graph, but rather identifies the set of structures compatible with the observed data and assumptions, it is best used to weed out statistically implausible causal relationships and hence generate testable causal hypotheses to be compared with theoretical expectations or followed up with targeted observations.
@@ 3.2 The FCIT algorithm applied to galaxies
For application to the NSA, we have the requirements that a method is 1) accurate in the presence of confounding latent variables, since it is highly unlikely that all relevant information is contained in the dataset, 2) able to accommodate highly non-linear functional relationships (see Fig. 1), and 3) efficient enough to analyse $N\approx\mathcal{O}(10^{5})$ objects in reasonable time. This is impossible for traditional algorithms, which scale as $N^{3}$ or worse when accounting for confounders and nonlinearity and become prohibitively memory intensive with $\gtrsim 500$ samples (Kalisch and Bühlmann, 2007; Strobl et al., 2018; Raghu et al., 2018).
To achieve this we adopt the newly-developed method Fast Causal Inference with Targeted Testing (FCIT; Ramsey et al. 2025a) as implemented in py-tetrad (Ramsey and Andrews, 2023). [^3] https://www.cmu.edu/dietrich/philosophy/tetrad/use-tetrad/tetrad-python.html, [^4] https://github.com/cmu-phil/py-tetrad This is a hybrid constraint-and-score-based algorithm designed to scale to high-dimensional settings while retaining robustness to latent confounding and nonlinear relationships.
FCIT proceeds in three main stages:
  1. 1.
    Skeleton discovery. Starting from a complete undirected graph, FCIT iteratively removes edges by testing conditional independences of increasing conditioning set size. Unlike the standard Fast Causal Inference algorithm Spirtes et al. (1993), FCIT restricts attention to conditioning sets that are local to the endpoints of each edge, which significantly reduces the number of tests required. It also incorporates discriminating path checks during edge removal, ensuring that edges are properly oriented before deciding on conditional independence.
  2. 2.
    Targeted blocking of remaining dependencies. For pairs of variables that remain dependent after standard local conditioning, FCIT applies a targeted testing strategy based on recursive blocking. This procedure searches for conditioning sets that block all remaining inducing paths between two variables by conditioning only on noncolliders. This targeted approach avoids the combinatorial explosion associated with exhaustive conditioning-set searches. Additional checks guarantee that interim and final graphs are well-formed PAGs when applied to data samples: if a candidate edge removal would violate this criterion, it is rejected. This is a departure from previous latent variable procedures that yield PAGs only theoretically.
  3. 3.
    Orientation and PAG construction. Once the adjacency structure is determined, FCIT applies orientation rules analogous to those used in FCI to identify unshielded colliders and propagate edge orientations, resulting in a PAG that encodes both causal directionality and ambiguity due to latent confounding.
By combining localised conditioning with targeted testing, FCIT achieves substantial computational savings relative to FCI while empirically retaining strong recovery of PAG structure in both simulated and real data settings. The resulting graphs are edge-minimal and correctly oriented, with an unprecedented runtime of only $\sim$1 minute for $4.5\times 10^{5}$ datapoints and 7 features. High accuracy on causal discovery benchmarks is demonstrated in Ramsey et al. (2025a).
Our independence test is use_basis_function_lrt. This implements a conditional independence test based on truncated basis-function expansions of the conditional mean functions. Specifically, to test for $X\perp Y\ |\ Z$ the variables $X$ and $Y$ are each expanded in an orthogonal basis (Legendre polynomials in our implementation), truncated at order given by the user-supplied hyperparameter truncation_limit. Conditional independence is then assessed by comparing a restricted model in which the basis coefficients of $X$ do not enter the regression for $Y$ to an unrestricted model in which they do using a likelihood ratio test, i.e. using the null and alternative hypotheses
$H_{0}:Y=f(Z)+\epsilon\;\;;\;\;\;\;H_{1}:Y=f(Z)+g(X)+\epsilon$ (1)
where $f$ and $g$ are basis-function expansions and $\epsilon$ a Gaussian noise. Independence is accepted if the improvement in log-likelihood allowing for the dependence on $X$ fails to exceed the $\chi^{2}$ threshold at the chosen significance level. This allows for flexible nonlinear dependencies while retaining a well-defined parametric test statistic.
For scoring we adopt use_basis_function_bic, which defines a likelihood by assuming additive noise models with finite variance after basis expansion. Specifically, the likelihood $\mathcal{L}$ is computed from the corresponding sums of squares of the residuals and therefore assumes that the noise term simply adds onto the dependent variable in each regression. Model complexity is penalised by the number of retained basis coefficients using the Bayesian information criterion (BIC; Schwarz 1978),
@@
where $N$ is the sample size used for the local regression and $k$ is the number of free regression coefficients in the basis‑function expansion for the child given its current parent set (Ramsey et al., 2025b). The BIC score balances goodness-of-fit against model complexity, with higher scores indicating preferred models in our convention. These assumptions are weaker than linear-Gaussian models while remaining computationally tractable at large sample sizes. The score-based search itself is done with the Best Order Score Search (BOSS; Andrews et al. 2023) algorithm.
This makes FCIT a hybrid algorithm: the constraint phase uses the basis‑function likelihood ratio test to prune edges, while the score phase (and the refinement and orientation phases) uses the basis‑function BIC. For the $p$-value threshold we use 0.01 in all cases, which we find not to affect the results appreciably. Note that FCIT is agnostic to the specific conditional independence test and score employed. We exploit this flexibility by combining FCIT with basis-function likelihood ratio tests and scores, allowing nonlinear dependencies to be detected without sacrificing scalability.
Two main hyperparameters affecting dataset-specific performance: penalty_discount and truncation_limit. The former controls how strongly graph complexity is penalised in the BIC score (Eq. 2). A higher value favours simpler graphs by removing more noise-dominated edges, at the expense of the quality of the conditional fit as described by $\mathcal{L}$. The latter controls the complexity of the local regression model through the number of polynomial basis terms included in the use_basis_function_lrt test as described above. Larger values allow more expressive models at the cost of runtime and enhanced BIC complexity penalty.